OpenStudy (anonymous):

Composite Function Check: Did i do this correctly? Given f(x)=x^3-8 g(x)=2x-4 find g(f(2)) This is what i did: 2^3-8=0 then i took the 0 and plugged it into g(x) so 2(0)-4=-4 this made my answer -4. Is that correct? Did i do this right??

OpenStudy (anonymous):

That's correct, though you could also do it symbolically...it'd be more work for the given problem though

OpenStudy (kc_kennylau):

That is correct

OpenStudy (anonymous):

ok awesome! More work or not i'll do what i gotta do to make sure i understand it and i'm doing it correctly :) Thanks so much!

OpenStudy (anonymous):

yep. It looks alright to me

OpenStudy (kc_kennylau):

but I would write the following: \[\hspace{14pt}g(f(2))\\=g(2^3-8)\\=g(0)\\=2(0)-4\\=-4\]

OpenStudy (anonymous):

Ok so can anyone check this one for me? Using the same function equations however g(f(0)) which i figured this way g(f(0))= g(0^3-8)= g=-8 2(-8)-4= -16-4 =-20

OpenStudy (anonymous):

That looks fine, but it wasnt' the question asked originally (Which awas g(f(2))

OpenStudy (anonymous):

nope it was a new question, i know i probably should have opened a new question but i didn't wanna seem repetitive